OpenStudy (anonymous):

If the surface area of a sphere is 144π cm2, what is the diameter?

OpenStudy (anonymous):

\[SA = 4 \pi r^2\] r = d/2

OpenStudy (anonymous):

So solve for r, then just multiply by 2, that's a pretty easy way.

OpenStudy (anonymous):

is the diameter 12?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

1/2 √144 = 6 6 x 2 = 12

OpenStudy (anonymous):

1/2?

OpenStudy (anonymous):

Where did you get that

OpenStudy (anonymous):

Somewhere on the internet, haha. I had to look all of this up because my teacher never taught me any of this..

OpenStudy (anonymous):

Do you know how to solve for r?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Have you not done algebra before?

OpenStudy (anonymous):

I did algebra last year but i'm an idiot so half of whatever I learned I ended up forgetting..

OpenStudy (anonymous):